Option #2: Personal Finance Application

Stock Valuation

  • Use the information provided to calculate the required rate of return on four stocks. Use both CAPM and the Dividend Discount Model. The market return is expected to be 8.5% and the risk-free rate is 2.0%.
Smith Company, Inc. Competitor 1 Competitor 2 Competitor 3
Beta 1.4 1.7 1.2 0.9
Dividend $2.00 $3.25 $2.00 $1.25
Price $26.00 $33.00 $45.00 $20.00
Growth 3.5% 3.5% 5.5% 1.5%
Answer #1

Smith Company Inc.
Using CAPM Required Rate =Risk free rate+Beta*(Expected Return -Risk free Rate) =2%+1.4*(8.5%-2%) =11.10%
Using Dividend Discount Rate =Dividend*(1+growth)/Price +Growth) =2*(1+3.5%)/26+3.5% =11.4615% or 11.46%

Competitor 1
Using CAPM Required Rate =Risk free rate+Beta*(Expected Return -Risk free Rate) =2%+1.7*(8.5%-2%) =13.05%
Using Dividend Discount Rate =Dividend*(1+growth)/Price +Growth) =3.25*(1+3.5%)/33+3.5% =13.69%

Competitor 2
Using CAPM Required Rate =Risk free rate+Beta*(Expected Return -Risk free Rate) =2%+1.2*(8.5%-2%) =9.80%
Using Dividend Discount Rate =Dividend*(1+growth)/Price +Growth) =2*(1+5.5%)/45+5.5% =10.19%

Competitor 3
Using CAPM Required Rate =Risk free rate+Beta*(Expected Return -Risk free Rate) =2%+0.9*(8.5%-2%) =7.85%
Using Dividend Discount Rate =Dividend*(1+growth)/Price +Growth) =1.25*(1+1.5%)/20+1.5% =7.84%